# A New Method to Simulate Dark Matter–Baryon Interactions and Application to an Isolated Disk Galaxy

### 25 Mar 2026 v1

@@ II.3 Baryon evolution
We next consider the updates to apply to the baryons. As noted above, we require that the updated baryon velocity distribution remain Maxwell-Boltzmann. This suggests an alternative approach in which we directly compute the rates-of-change of the parameters of the Maxwell-Boltzmann distribution: the bulk velocity and the temperature. These are
@@
These can be evaluated using techniques very similar to those described by Ali-Haïmoud (2019), augmented by an integral over space as in the dark matter case. Details are provided in Appendix B. The result is
@@
where $g_{ij}$ is the overlap integral as before and $\mathcal{A}$ and $\mathcal{B}$ are defined
@@
Notice the appearance of the momentum transfer cross section, $\overline{\sigma}$. If the momentum transfer cross section has a power law dependence on $v$, then these integrals $\mathcal{A}$ and $\mathcal{B}$ have simple forms in terms of special functions, given in Appendix B. In simulation, these rates-of-change can simply be treated as new sources of acceleration and heating, added onto the total acceleration and heating rates which are already accumulated for each particle and integrated by the preferred scheme of the simulation.
These rates-of-change are statistically equal-and-opposite to the Monte Carlo sampling procedure we have outlined for the dark matter. Momentum and energy are conserved in expectation value, which we prove in Appendix C.
We note that the dark matter and baryon calculations end up using different integrals over the differential cross section: the dark matter scattering probability depends on the total cross section $\sigma$, while the gas momentum and heat exchange rates depend on the momentum transfer cross section $\overline{\sigma}$. In the self-interacting dark matter literature, it is often assumed that one can approximate a non-isotropic cross section in the scattering probability by using isotropic scatterings with the momentum transfer cross section (Tulin and Yu, 2018). For dark matter–baryon interactions, such an assumption must be carefully considered, as it could break the equivalence between the two calculations.
@@ II.4 GIZMO implementation
The core of the implementation is a nearest-neighbors calculation, where dark matter particles seek gas particles and vice versa. Each particle searches in a radius equal to twice its softening length. With this search radius, each pair of particles which overlaps is guaranteed to be found by the neighbor with larger softening length, but may or may not be found by the smaller neighbor. As such, we must search in both directions (dark matter $\rightarrow$ gas and vice versa); to ensure that each pair is processed only once, we skip a pair whenever the searching particle has the smaller softening length.
To process pair $(i,j)$ (taking $i$ to be dark matter, $j$ gas), we first compute the overlap integral $g_{ij}$. We then compute the contribution of particle $i$ to the acceleration ${\rm d}\bm{V}_{j}/{\rm d}t$ and heating rate ${\rm d}T_{j}/{\rm d}t$ of particle $j$, which are accumulated over all pairs that include $j$. Next, we draw a random sample from the gas velocity distribution $\mathcal{G}(\bm{v}-\bm{V}_{j},T_{j}/\mu_{j})$ and use it to compute the scattering probability $P_{ij}$. If the computed scattering probability $P_{ij}$ is too large, we reduce the timestep size for the dark matter particle for the next timestep. The specific criterion we use is
@@
which requires smaller timesteps if $m_{\chi}\lesssim\mu_{j}$, because in that case scatters are expected to impart a large change on the dark matter velocity. We draw a random number $r$ uniformly between 0 and 1 to decide whether to scatter the dark matter particle. If we choose to scatter (if $r<P_{ij}$), we choose a new direction $\hat{\bm{v}}_{\chi B}^{\prime}$ isotropically, then compute $\bm{v}_{\chi}^{\prime}$ from Equation 1 and apply it to particle $i$.
@@
where $(\cdot)_{k}$ is the Pochhammer symbol. In our implementation, we switch to this expansion whenever $|x|\geq 50$ (using the GSL implementation otherwise), and we cut off the sum at five terms. This gives fractional errors of less than $10^{-14}$ for all $x$ of interest when compared against the Boost Math library implementation. [^3] See the Boost library’s excellent discussion of the methods they employ to calculate this function accurately: https://www.boost.org/doc/libs/latest/libs/math/doc/html/math_toolkit/hypergeometric/hypergeometric_1f1.html.
@@ II.5 Comparison to existing methods
The method we use to treat dark matter bears many similarities to methods for simulating self-interacting dark matter (e.g., Rocha et al., 2013). Our treatment of baryons borrows much technology from analyses of dark matter–baryon interactions in linear cosmology (e.g., Dvorkin et al., 2014; Ali-Haïmoud, 2019). At the time of publication of this article, however, the only existing method specifically designed for simulating dark matter–baryon interactions is Fischer et al. (2025), hereafter F25. We give in this section a detailed comparison of our method to the F25 scheme for “rare, large-angle scatters”.
@@
The key difference is our treatment of baryons. We treat baryons by using the requirement that their velocity distribution remain Maxwell-Boltzmann at all times to calculate moments of the Boltzmann equation, obtaining momentum and heat transfer rates that can be directly applied to the gas macroparticles. F25 instead develop a scheme to apply the probabilistic scatters to the baryons. The issue they must solve, and which we sidestep, is how to conduct a Monte Carlo scattering between two different particle species when the macroparticles represent different numbers of underlying particles: specifically, when the dark matter weight $M_{\chi}/m_{\chi}$ differs from the baryon weight $M_{B}/m_{B}$. They do so by constructing a “virtual” baryon particle with mass $M_{\rm virt}$ set so that $M_{\rm virt}/m_{B}=M_{\chi}/m_{\chi}$, giving both macroparticles equal weight and making scattering straightforward. After scattering, the changes in momentum and energy of the virtual particle are distributed back to the simulation baryon particle, guaranteeing conservation of energy and momentum on a per-scattering basis.
The difficulty with this method is ensuring that the internal energy of the simulation particle remains positive after applying the change in energy of the virtual particle. In particular, to ensure that scatters with all possible sampled baryon velocities $\bm{v}_{B}$ and scattering angles $\Omega$ leave the baryon macroparticle with positive energy, one needs the virtual baryon mass $M_{\rm virt}$ to be only a fraction of the baryon macroparticle mass $M_{B}$. For a given ratio of true particle masses, $m_{\chi}/m_{B}$, this sets an upper limit on the dark matter macroparticle mass $M_{\chi}$ and, when the dark matter mass is small, $m_{\chi}\lesssim m_{B}$, can mean requiring that $M_{\chi}$ be much smaller than $M_{B}$. Such a requirement would be computationally infeasible in many cosmological and galaxy simulations, where the dark matter is the dominant component of the simulation by mass and where high resolution for the baryons is needed. If this mass ratio is not satisfied, F25 note that one can simply reject the fraction of scatters which would lead to negative internal energies. However, they also note that, even though these are a small fraction of all possible scatters, they yield the largest momentum and energy exchange, so cutting them could lead to an underestimate of the effects of dark matter–baryon interactions.
Our method, by divorcing the treatment of the baryons from that of the dark matter, necessarily gives up conservation of energy and momentum on a per-scattering basis, although we prove conservation in the expectation values. However, this split treatment does not impose any restrictions on the masses of the macroparticles, which makes accurate simulation feasible for low mass dark matter.
There is one other, smaller difference which comes about due to our differing treatment of the baryons. The random kicks received by the baryon macroparticles in F25 are applied to the particle velocities $\bm{V}_{j}$, which the particles use as estimates of the local bulk velocity. Since the kicks are randomly oriented, they add variance to the bulk velocity field, effectively siphoning energy out of the internal energies of the particles and into their velocities. In order to prevent a drastic under-reporting of the internal energies of the particles, F25 use viscosity in the baryons to ensure that these random bulk motions are absorbed into the internal energies. To contrast, our approach requires no consideration of viscosity, because the momentum and energy transfer is already split into bulk and thermal components.
@@ III Tests
We have verified our implementation by constructing a test problem that isolates the dark matter–baryon interactions. Explicit details and results are given in Appendix E; we present a brief summary here. The test problem consists of a box with periodic boundary conditions filled uniformly with dark matter and gas of specified densities. The gas particles start at rest and are given some initial internal energy. The dark matter particles are given velocities sampled from a Maxwell-Boltzmann distribution with some bulk velocity and temperature. All physics modules in GIZMO are disabled except for dark matter–baryon interactions. We test a wide variety of initial conditions and mass resolutions for our fiducial dark matter model, which has mass $m_{\chi}=2\,m_{p}$ and velocity-independent cross section $\sigma=10\ {\rm mb}$ ($10^{-26}\ {\rm cm}^{2}$). We also test alternative dark matter models for a limited set of initial conditions, including different dark matter masses, $m_{\chi}\in\{10\,\text{GeV},100\,\text{MeV},10\,\text{MeV}\}$, and velocity-dependent cross sections, $n\in\{-2,-4\}$.
The test problem is designed to facilitate comparison to a direct numerical integration of a system of differential equations derivable under the assumption that the dark matter is Maxwell-Boltzmann distributed at all times. Of course, this assumption does not and should not be expected to hold, as the dark matter departs from equilibrium while transferring energy and momentum to the gas. However, for many choices of the test setup, we find that the dark matter velocity distribution does not stray too far from a thermal distribution, and so its behavior agrees well with the Maxwell-Boltzmann calculation. For other setups where the dark matter evolution departs substantially from a thermal distribution, we cannot directly predict the evolution of the system, but we still verify that the two species equilibrate sensibly and that momentum and energy are conserved. A detailed discussion of our test choices, the comparison to the naïve Maxwell-Boltzmann prediction, and the results of our simulations is given in Appendix E.
Coincidentally, our test setup is almost identical to that which was employed in Fischer et al. (2025). As such, we re-create the most analogous of their tests (which is detailed in their Section 3.4). Our code produces strikingly similar behavior to what they report.
@@ IV Galaxy simulations
Figure 1: Top-down view of the gas in the center of the disk (left) and the entire stellar disk (right) at a variety of cross sections (increasing from top to bottom) and times (increasing from left to right). The largest two cross sections show the development of an extremely compact and dense object in the central gas, as well as the concentration of stars in the center. No significant differences are seen between CDM and the smaller cross sections.
@@
With these initial conditions, we perform a suite of simulations using our modified version of GIZMO (Hopkins, 2015) to include dark matter–baryon interactions. We choose a dark matter particle mass of $m_{\chi}=2\,m_{p}$ and velocity-independent cross sections ($n=0$) of $\{0,10,30,100,300,1000\}\ \text{mb}$. We match the parameters and physics used in Kim et al. (2016), including radiative cooling via the Grackle library (Smith et al., 2017), metal production, star formation, a simple thermal stellar feedback model, a Jeans pressure floor (Robertson and Kravtsov, 2008), and gravitational softening lengths of 224 pc for all particle types. For the kernel used, this choice of softening length means gravity is exactly Newtonian at distances greater than 80 pc.
Additionally, we use adaptive softening lengths for the gas. For dark matter, we follow the SIDM implementation in GIZMO (Rocha et al., 2013), which uses adaptive softening lengths only for determining which particles overlap and for the calculation of the overlap integral $g_{ij}$. Importantly, the gravity solver uses fixed softening lengths for dark matter. In preliminary tests, we saw little change when using fixed dark matter softening lengths for the dark matter–baryon interactions. Adaptive softening lengths are initialized to 224 pc, and can vary down to a minimum of 20% of this value (44.8 pc).
We evolve the simulations for approximately 1 Gyr. Since we do not attempt to simulate the cosmic baryon cycle, the gas in the galaxy gets used up for star formation, shutting off the dark matter–baryon interactions prematurely and therefore setting a limit on the length of the simulation. While longer simulations would be interesting, especially for smaller cross sections where the impact of interactions takes longer to build up, these would require a model for gas accretion during the evolution of the galaxy.

### 17 Apr 2026 v2

@@ II.3 Baryon evolution
We next consider the updates to apply to the baryons. The updated baryon velocity distribution is required to remain Maxwell-Boltzmann. This suggests an alternative approach in which we directly compute the rates-of-change of the parameters of the Maxwell-Boltzmann distribution: the bulk velocity and the temperature. These are
@@
These can be evaluated using techniques very similar to those described by Ali-Haïmoud (2019), augmented by an integral over space (like in the dark matter case). Details are provided in Appendix B. The result is
@@
where $g_{ij}$ is the overlap integral as before and $\mathcal{A}$ and $\mathcal{B}$ are defined:
@@
Notice the appearance of the momentum transfer cross section, $\overline{\sigma}$. If the momentum transfer cross section has a power-law dependence on $v$, then these integrals $\mathcal{A}$ and $\mathcal{B}$ have simple forms in terms of special functions, given in Appendix B. In simulation, these rates-of-change can be treated directly as new sources of acceleration and heating, added onto the total acceleration and heating rates, which are already accumulated for each particle and integrated by the preferred scheme of the simulation.
These rates-of-change are statistically equal and opposite to the Monte Carlo sampling procedure we have outlined for the dark matter. Momentum and energy are conserved in expectation value, which we prove in Appendix C.
We note that the dark matter and baryon calculations end up using different integrals over the differential cross section: the dark matter scattering probability depends on the total cross section $\sigma$, while the gas exchange rates depend on the momentum transfer cross section $\overline{\sigma}$. In the self-interacting dark matter literature, it is often assumed that one can approximate a non-isotropic cross section in the scattering probability by using isotropic scatterings with the momentum transfer cross section (Tulin and Yu, 2018). For dark matter–baryon interactions, such an assumption must be carefully considered, as it could break the equivalence between the two calculations.
@@ II.4 GIZMO implementation
The core of the implementation is a nearest-neighbors calculation, where dark matter particles seek gas particles and vice versa, during the “kick” step of the GIZMO leapfrog integrator. Each particle searches in a radius equal to twice its softening length. With this search radius, each pair of particles that overlaps is guaranteed to be found by the neighbor with larger softening length, but may or may not be found by the smaller neighbor. Accordingly, we must search in both directions (dark matter $\rightarrow$ gas and vice versa); to ensure that each pair is processed only once, we skip a pair whenever the searching particle has the smaller softening length.
To process pair $(i,j)$ (taking $i$ to be dark matter, $j$ gas), we first compute the overlap integral $g_{ij}$. We then compute the contribution of particle $i$ to the acceleration ${\rm d}\bm{V}_{j}/{\rm d}t$ and heating rate ${\rm d}T_{j}/{\rm d}t$ of particle $j$, which are accumulated over all pairs that include $j$. Next, we draw a random sample from the gas velocity distribution $\mathcal{G}(\bm{v}-\bm{V}_{j},T_{j}/\mu_{j})$ and use it to compute the scattering probability $P_{ij}$. If the scattering probability $P_{ij}$ is too large, we reduce the timestep size for the dark matter particle for the next timestep. The criterion we use is
@@
which requires smaller timesteps if $m_{\chi}\lesssim\mu_{j}$ (because in that case scatters are expected to impart a large change on the dark matter velocity). [^6] Note that the sum $\sum_{j}P_{ij}$ should also remain small. For $m_{\chi}=2\,m_{p}$, Equation 19 imposes a maximum of $P_{ij}\lesssim 10^{-3}$, with most probabilities significantly less than this threshold, so this is safely satisfied in our simulations. We draw a random number $r$ uniformly between 0 and 1 to decide whether to scatter the dark matter particle. If we choose to scatter (if $r<P_{ij}$), we choose a new direction $\hat{\bm{v}}_{\chi B}^{\prime}$ isotropically, then compute $\bm{v}_{\chi}^{\prime}$ from Equation 1 and apply it to particle $i$.
@@
where $(\cdot)_{k}$ is the Pochhammer symbol. In our implementation, we switch to this expansion whenever $|x|\geq 50$ (using the GSL implementation otherwise), truncating the sum at five terms. This gives fractional errors of less than $10^{-14}$ for all $x$ of interest when compared against the Boost Math library implementation. [^7] See the Boost library’s excellent discussion of the methods they employ to calculate this function accurately: https://www.boost.org/doc/libs/latest/libs/math/doc/html/math_toolkit/hypergeometric/hypergeometric_1f1.html.
@@
Lastly, we note that our implementation properly accounts for the conversion from comoving to physical coordinates and the Hubble flow in GIZMO, although this is not directly tested in the simulations in this paper.
@@ II.5 Comparison to existing methods
The method we use for dark matter is very similar to methods for simulating self-interacting dark matter (e.g., Rocha et al., 2013). Our treatment of baryons borrows much technology from analyses of dark matter–baryon interactions in linear cosmology (e.g., Dvorkin et al., 2014; Ali-Haïmoud, 2019). At the time of publication of this article, however, the only existing method specifically designed for simulating dark matter–baryon interactions is that of Fischer et al. (2025) (F25). We give in this section a detailed comparison of our method to the F25 scheme for “rare, large-angle scatters”; F25 also gives a scheme for “frequent, small-angle scatters” specialized to a certain class of cross section, which largely avoids the pitfalls we describe but is not analogous to our method.
@@
The key difference is our handling of baryons. We use the fact that the baryon velocity distribution remains Maxwell-Boltzmann at all times to calculate moments of the Boltzmann equation. This gives us momentum and heat transfer rates that can be directly applied to the gas macroparticles. F25 instead develops a scheme to apply the probabilistic scatters to the baryons. The issues they must solve, and which we sidestep, are how to conduct a scattering between macroparticles that represent (1) different numbers of physical particles (i.e., when $M_{\chi}/m_{\chi}\neq M_{B}/m_{B}$), and (2) different kinds of velocity distributions (i.e., Dirac delta for dark matter and Maxwell-Boltzmann for baryons). F25 handles these issues by constructing a “virtual” baryon particle for scattering. The virtual particle has mass $M_{\rm virt}$ set so that $M_{\rm virt}/m_{B}=M_{\chi}/m_{\chi}$, which makes it straightforward to apply a scattering. After scattering, the changes in momentum and energy of the virtual particle are distributed back to the simulation baryon particle, which makes use of the Maxwell-Boltzmann distribution nature of baryon macroparticles.
The virtual baryon approach has a few key advantages. One is that it is marginally more efficient: baryons are only updated when a partner dark matter particle is updated, and these updates do not require the calculation of any special functions. Another is that energy and momentum are conserved on a per-scattering basis. We contrast these against our code, where baryons receive small, smooth updates every timestep, these updates require calculating special functions, and the split approach means we can only conserve energy and momentum on average.
The primary difficulty with the virtual baryon approach is ensuring that the internal energy of the simulation particle remains positive after the scattering. This is enforced by rejecting scatterings whenever they result in a negative internal energy. The rate of these rejections primarily depends on how the mass ratio of the macroparticles, $M_{\chi}/M_{B}$, compares to that of the true particles, $m_{\chi}/m_{B}$. If a large fraction of scatters are rejected, the energy transfer between species will be biased.
We have performed a numerical experiment using the F25 algorithm to determine what macroparticle mass ratios are valid for a range of true particle mass ratios. A detailed discussion of our numerical experiment is given in Appendix E. The rate at which scatters are rejected and the resulting bias of the overall energy transfer are shown in Figure 1. The vertical bar denotes the upper boundary of the “safe zone” given in Equation 12 of F25; one sees that their safe zone is robust and even conservative by a factor of 2, depending on the accuracy required. In general, the accuracy of the F25 virtual baryon approach rapidly and severely degrades whenever $M_{\chi}/M_{B}\gtrsim\frac{1}{3}\,m_{\chi}/m_{B}$.
Figure 1: The rejection rate (top) and fractional error of the energy transfer (bottom) for the Fischer et al. (2025) (F25) method as a function of the macroparticle mass ratio $M_{\chi}/M_{B}$ divided by the physical particle mass ratio $m_{\chi}/m_{B}$. The region that is expected to be “safe” according to Equation 12 of F25 is marked with a vertical line and a leftward arrow. In agreement with this expectation, we find that the algorithm is valid whenever $M_{\chi}/M_{B}\ll m_{\chi}/m_{B}$, as the rejection rate is negligible and the energy transfer matches the expectation. For $M_{\chi}/M_{B}\gtrsim m_{\chi}/m_{B}$, the transition to rejecting nearly all scatters happens rather abruptly, leading to a severe underestimate of the energy transfer rate.
For small dark matter particle masses, $m_{\chi}\lesssim m_{B}$, restricting to the safe zone means that the dark matter macroparticles need much lower mass (i.e., higher resolution) than baryons. Such a requirement is computationally infeasible in many cosmological and galaxy simulations, where the dark matter is the dominant component of the simulation by mass and high resolution for the baryons is needed. Because our method makes different trade-offs, it remains accurate even in the $M_{\chi}/M_{B}\gtrsim m_{\chi}/m_{B}$ regime.
In summary, the F25 approach has computational advantages, including exact conservation of momentum and energy, which in some regimes make it more efficient than the method we have developed. However, F25 is only accurate when the $M_{\chi}/M_{B}\lesssim\tfrac{1}{3}\,m_{\chi}/m_{B}$. For small dark matter masses $m_{\chi}\lesssim m_{B}$, this restriction would make many cosmological simulations infeasible. However, our method remains accurate without this restriction, enabling much lower computational cost. Further, as will be described below, we find nearly identical performance to the F25 method in a test problem where the F25 method is accurate.
@@ III Tests
We have verified our implementation by constructing a test problem that isolates the dark matter–baryon interactions. Explicit details and results are given in Appendix F; we present a brief summary here. The test problem consists of a box with periodic boundary conditions filled uniformly with dark matter and gas of specified densities. The gas particles start at rest and are given some initial internal energy. The dark matter particles are given velocities sampled from a Maxwell-Boltzmann distribution with some bulk velocity and effective temperature. All physics modules in GIZMO are disabled except for dark matter–baryon interactions. We test a wide variety of initial conditions and mass resolutions for our fiducial dark matter model, which has mass $m_{\chi}=2\,m_{p}$ and velocity-independent cross section $\sigma=10\ {\rm mb}$ ($10^{-26}\ {\rm cm}^{2}$). We also test alternative dark matter models for a limited set of initial conditions, including different dark matter masses, $m_{\chi}\in\{10\,\text{GeV},100\,\text{MeV},10\,\text{MeV}\}$, and velocity-dependent cross sections, $n\in\{-2,-4\}$.
The test problem is designed to facilitate comparison to a direct numerical integration of a system of differential equations derivable under the assumption that the dark matter is Maxwell-Boltzmann distributed at all times. This assumption does not and should not be expected to hold, as the dark matter departs from equilibrium while transferring energy and momentum to the gas. However, for many choices of the test setup, we find that the dark matter velocity distribution does not stray too far from a thermal distribution, and so its behavior agrees well with the Maxwell-Boltzmann calculation. For other setups where the dark matter evolution departs substantially from a thermal distribution, we cannot directly predict the evolution of the system, but we still verify that the two species equilibrate sensibly and that momentum and energy are conserved. A detailed discussion of our test choices, the comparison to the naïve Maxwell-Boltzmann prediction, and the results of our simulations are given in Appendix F.
Coincidentally, our test setup is almost identical to the one employed in Fischer et al. (2025), so we re-create the most analogous of their tests (detailed in their Section 3.4). Our code produces essentially indistinguishable behavior to what they report.
@@ IV Galaxy simulations
Figure 2: Top-down view of the gas in the center of the disk (left) and the entire stellar disk (right) at a variety of cross sections (increasing from top to bottom) and times (increasing from left to right). The two largest cross sections show the development of an extremely compact and dense object in the central gas, as well as the concentration of stars in the center. No significant differences are seen between CDM and the smaller cross sections.
@@
With these initial conditions, we perform a suite of simulations using our modified version of GIZMO (Hopkins, 2015) to include dark matter–baryon interactions. We choose a dark matter particle mass of $m_{\chi}=2\,m_{p}$ and velocity-independent cross sections ($n=0$) of $\{0,10,30,100,300,1000\}\ \text{mb}$. We match the parameters and physics used in Kim et al. (2016), including radiative cooling via the Grackle library (Smith et al., 2017), metal production, star formation, a simple thermal stellar feedback model, a Jeans pressure floor (Robertson and Kravtsov, 2008), and gravitational softening lengths of 224 pc for all particle types. For the kernel used, gravity is exactly Newtonian at distances greater than $1/2.8$ times the softening length (80 pc in this case).
Additionally, we use adaptive softening lengths for the gas. For dark matter, we follow the SIDM implementation in GIZMO (Rocha et al., 2013), which uses adaptive softening lengths only for determining which particles overlap and calculating the overlap integral $g_{ij}$. These adaptive softening lengths are initialized to 224 pc, and can vary down to a minimum of 20% of this value (44.8 pc). Importantly, the gravity solver uses fixed softening lengths for dark matter. In preliminary tests, we saw little change when instead using fixed dark matter softening lengths to compute the overlap integral.
We evolve the simulations for approximately 1 Gyr. Since we do not attempt to simulate the cosmic baryon cycle, some of the gas in the galaxy is consumed by star formation, which prematurely slows the dark matter–baryon interactions and therefore sets a limit on the length of the simulation. While longer simulations would be interesting, especially for smaller cross sections where the impact of interactions takes longer to build up, these would require a model for gas accretion during the evolution of the galaxy.